6. (a) For normal distribution with mean zero and variance 2

σ show that:


2

(| |) σ

π

E x =

 (


1
Expert's answer
2021-03-25T06:45:49-0400

Given,

Mean "\mu =\ 0\\"

Variance "= {\sigma}^2"

To prove: "\sigma=\sqrt{E(X^2)}"

We know that

"Var(X)= E(X^2)-E(X)^2"

"Var(X)=E(X^2)-\mu^2"

But "\mu = 0"

So, "Var(X)= E(X^2)\\\Rightarrow \sigma^2 = E(X^2)"

Hence, "\sigma = \sqrt{E(X^2)}"